OpenStudy (nuccioreggie):

Please show every single step. x + 2 = 12 x - 9 = 26 5x = 40 x/4 = 3 2x - 4 = -2 2x + 7 = 19 (x/3) + 4 = 10 3x + 1 = 25 3x - 2 < 13 2x + 7 > 21

OpenStudy (anonymous):

For the first one....

OpenStudy (greatlife44):

x + 2 = 12 for the first one what do you think.

OpenStudy (anonymous):

x + 2 = 12 You need to undo the addition of 2 so that x is by itself. Understand? Subtract x by both sides. x + 2 - 2 = 12 - 2 x = 12 - 2 What is 12 - 2?

OpenStudy (nuccioreggie):

I dont know i need help with them tho

OpenStudy (anonymous):

Look at what I said

OpenStudy (anonymous):

Sorry in my response i meant subtract TWO by both sides not x

OpenStudy (nuccioreggie):

10

OpenStudy (greatlife44):

You isolate need to get x by itself. if you see addition you do subtraction and whatever you do to one side you do to the other

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

the answer for the first one is 10

OpenStudy (greatlife44):

@nuccioreggie you know your answer is correct because 10+2 = 12

OpenStudy (anonymous):

Do you know how to do number 2?

OpenStudy (anonymous):

remember that the 9 is being subtracted, so do the opposite of subtraction to both sides. the opposite of subtraction is....?

OpenStudy (nuccioreggie):

x + 2 = 12 You need to undo the addition of 2 so that x is by itself. Understand? Subtract x by both sides. x + 2 - 2 = 12 - 2 x = 12 - 2 What is 12 - 2=10 So thats the answer for nummber 1

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

except the x should be 2 my mistake

OpenStudy (nuccioreggie):

Which x

OpenStudy (anonymous):

You need to undo the addition of 2 so that x is by itself. Understand? Subtract TWO (not x) by both sides.

OpenStudy (anonymous):

@nuccioreggie

OpenStudy (nuccioreggie):

x + 2 = 12 You need to undo the addition of 2 so that x is by itself. Understand? Subtract x by both sides. x + 2 - 2 = 12 - 2 2 = 12 - 2 What is 12 - 2=10 @DIDITHEDAD

OpenStudy (anonymous):

The "Subtract x by both sides" that needs to be "Subtract 2 by both sides"

OpenStudy (anonymous):

x + 2 = 12 You need to undo the addition of 2 so that x is by itself. Understand? Subtract 2 by both sides. x + 2 - 2 = 12 - 2 x = 12 - 2 x = 10 (Copy and paste that)

OpenStudy (nuccioreggie):

Wait to can you just re write number one out for me please

OpenStudy (nuccioreggie):

Lol nvm

OpenStudy (anonymous):

x - 9 = 26 What do you think we do?

OpenStudy (nuccioreggie):

x + 2 = 12 You need to undo the addition of 2 so that x is by itself. Understand? Subtract 2 by both sides. x + 2 - 2 = 12 - 2 x = 12 - 2 x = 10 So this the answer for number one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What do we do for number two? x - 9 = 26 It is being subtracted, so we do the opposite....the opposite of subtraction is?

OpenStudy (nuccioreggie):

Ok so for number 2

OpenStudy (anonymous):

Opposite of subtraction?

OpenStudy (nuccioreggie):

Addition

OpenStudy (anonymous):

Yes...so x - 9 = 26 Add 9 to both sides x - 9 + 9 = 26 + 9 x = 26 + 9 what is 26 + 9?

OpenStudy (nuccioreggie):

35

OpenStudy (anonymous):

Yup!

OpenStudy (nuccioreggie):

x - 9 = 26 Add 9 to both sides x - 9 + 9 = 26 + 9 x = 26 + 9 what is 26 + 9= 35 So thats the answer for number 2?

OpenStudy (anonymous):

5x = 40 5x is 5 times x. So this is multiplication... Opposite of multiplication?

OpenStudy (anonymous):

For number two, x - 9 = 26 Add 9 to both sides x - 9 + 9 = 26 + 9 x = 35

OpenStudy (nuccioreggie):

Ok 1sec im writing these down ok so i might take a little time to respond back but just dont leave the question im just writing these down ok

OpenStudy (anonymous):

OKay

OpenStudy (nuccioreggie):

Ok im ready for 3 now

OpenStudy (anonymous):

5x = 40 5x is 5 times x... times is multiplication... opposite of multiplication is what?

OpenStudy (nuccioreggie):

Division And i didnt put that my computer does that sumtimes

OpenStudy (anonymous):

Yup. So divide 5 by both sides 5x = 40 5x/5 = 40/5 (The / means divided by) x = 40/5 What is 40 divided by 5? 

OpenStudy (nuccioreggie):

8

OpenStudy (anonymous):

x = 8, CORRECT

OpenStudy (anonymous):

5x = 40 Divide both sides by 5... 5x/5 = 40/5 x = 8

OpenStudy (anonymous):

x/4 = 3 x divided by 4 Opposite of division is what?

OpenStudy (nuccioreggie):

Ok 4

OpenStudy (anonymous):

What is the opposite of division?

OpenStudy (nuccioreggie):

Multiplication

OpenStudy (anonymous):

Soooooo do this equation by yourself! x/4 = 3 Multiply both sides by 4

OpenStudy (nuccioreggie):

64?

OpenStudy (anonymous):

No...Try again...